OpenStudy (anonymous):

can (5n^3)^2 + 2n^2(n^3+4n^4) be simplified further than 25n^6 + 2n^2(n^3+4n^4). All of my answer checks say no. Just verifying my answer. Thank you!

OpenStudy (anonymous):

i would be 25n^6 + 2n^5*(1+4n)

OpenStudy (raden):

i think u need to expand it for (5n^3)^2 = 25n^6 and expand of 2n^2(n^3+4n^4) = 2n^5 + 8n^6 just add of them be: 25n^6 + 2n^5 + 8n^6 = 33n^6 + 2n^5

OpenStudy (anonymous):

ed, your answer checks out but I have zero idea how you arrived to the conclusion. RadEn, Thank you! That completely makes sense to add the like terms in the end.

OpenStudy (raden):

you're welcome :)